A 20.64 g ice cube at -1.93 oC is place in an aluminum cup whose initial temperature is 137.9 oC. The system comes to an equilibrium temperature of 17.96 oC. What is the mass in grams of the cup? The specific heat of ice is 2.093 J/g oC and that of aluminum is 0.901 J/g oC.
Answer #1

heat gained by ice Q1 = M_ice*dT1 + Mice*Lf +Mice*Cw*dT2

Q1 = (20.64*2.093*1.93)+(20.64*334)+(20.64*4.19*17.96)


Q1 = 8530.345 J


heat loss by cup = M_cup*Cal*dT

Q2 = Mcup*0.901*(1379-17.96) = 1226.29704*M_cup

from calorimetry

Q2 = Q1


1226.29704*M_cup = 8530.345

Mcup = 6.956 g
Answer #2

use the equation q=(m*C*deltaT),

where m equals the mass,

C equals the specific heat,

deltaT equals the change in temperature, and

q equals the heat